OpenStudy (anonymous):

4/5*3/2+1/3*1/2 helpp. i know its basic lol

OpenStudy (anonymous):

convert the equations to the LCM

OpenStudy (saifoo.khan):

BoDMAS.

OpenStudy (aravindg):

saif help meeeeeeeeeeeeeee

OpenStudy (mimi_x3):

Use a calculator mans , common sense (: 41/30

OpenStudy (anonymous):

my calculator is missing lol

OpenStudy (saifoo.khan):

thiss is impossible. Soo Hard

OpenStudy (anonymous):

lol ok thank youuu

OpenStudy (mimi_x3):

The "legend" is saying that it is hard :O

OpenStudy (saifoo.khan):

akshay your one is ULRTA HARD/

OpenStudy (mimi_x3):

akshay?! he's not even here

OpenStudy (anonymous):

i was gonna say.. lol

OpenStudy (saifoo.khan):

he will be.

OpenStudy (anonymous):

/30

OpenStudy (saifoo.khan):

wow.

OpenStudy (saifoo.khan):

nothing/ 30

OpenStudy (saifoo.khan):

lol

OpenStudy (mimi_x3):

lols , how can it be 0 >_<

OpenStudy (anonymous):

lol does anoyone wanna be awesome and show the work. im being extra lazyyy lol

OpenStudy (saifoo.khan):

im being Ultra Lazy now.

OpenStudy (anonymous):

lol i just hate online homework. i get distracted super easily. with facebook and twitter and tumblr open. and everything else ahha

OpenStudy (mimi_x3):

It's quite obvious , just use a calculator mans . It's pointless doing it by hand

OpenStudy (saifoo.khan):

lol. are u in online college?

OpenStudy (anonymous):

no im in regular college. but the homework is online. i had to take a basic algebra class cause i suck at math lol

OpenStudy (saifoo.khan):

Oh Right, cool.

OpenStudy (anonymous):

4/5=24/30 3/2=15/30 1/3=10/30 1/2=15/30 24/30*15/30+10/30*15/30 360/30+150/30 510/30

OpenStudy (mimi_x3):

lols , your actually bothered , mrbone xD

OpenStudy (anonymous):

the answer was 40/30. or something like that lol

OpenStudy (anonymous):

not 510

OpenStudy (anonymous):

what did i do wrong then?

OpenStudy (anonymous):

idk lol

OpenStudy (saifoo.khan):

\[\Huge \frac{41}{30}\]

OpenStudy (mimi_x3):

According to my calculator it says 41/30.

OpenStudy (anonymous):

yeah. the website was all "good f'ing job" when i typed that in lol